OpenStudy (sphott51):

Which equations gives the amount P that should be invested at 8% per annum , compounded annually, so that $2000 will be available after 7 years? A. P=2000*7^(1.08) b. P=2000(0.08)^7 c. P=2000(1.08)^7 d. P=2000/(0.08)^7 e. P=2000/(1.08)^7

jimthompson5910 (jim_thompson5910):

The formula you use is A = P*(1+r/n)^(n*t) where, A = final amount P = initial amount r = interest rate in decimal form n = number of times you compound the money per year t = number of years

OpenStudy (sphott51):

so it's P= 2000/(0.08)^7

jimthompson5910 (jim_thompson5910):

`Which equations gives the amount P that should be invested at 8% per annum , compounded annually, so that $2000 will be available after 7 years? ` that info gives A = 2000 P = unknown r = 0.08 n = 1 t = 7

jimthompson5910 (jim_thompson5910):

`so it's P= 2000/(0.08)^7` incorrect

OpenStudy (sphott51):

Idk then

jimthompson5910 (jim_thompson5910):

you were really close but you made a mistake with the 1+r/n part. Try that again

OpenStudy (sphott51):

I know its not b or d.. Maybe maybe its c?

jimthompson5910 (jim_thompson5910):

sounds like you're randomly guessing at this point

OpenStudy (sphott51):

Sorry I meant a, cause 8% can't be 1.08.. Right..?

jimthompson5910 (jim_thompson5910):

notice how 1+r/n = 1+0.08/1 = 1.08

jimthompson5910 (jim_thompson5910):

another hint: solve for P in A = P*(1+r/n)^(n*t) so you'll have \[\LARGE P = \frac{A}{(1+r/n)^{n*t}}\]

OpenStudy (sphott51):

Oh so.. E then..?

jimthompson5910 (jim_thompson5910):

A = 2000 P = unknown r = 0.08 n = 1 t = 7 \[\LARGE P = \frac{A}{(1+r/n)^{n*t}}\] \[\LARGE P = \frac{2000}{(1+0.08/1)^{1*7}}\] \[\LARGE P = \frac{2000}{1.08^{7}}\]

jimthompson5910 (jim_thompson5910):

yeah E